# DeepEN: Personalized Enteral Nutrition for Critically Ill Patients using Deep Reinforcement Learning

### v1

@@ 3.2.2 Baselines
    Expert Guidelines (EG) Policy: This rule-based policy encodes clinical recommendations drawn from the American Society for Parenteral and Enteral Nutrition (ASPEN) guidelines [3], in combination with more recent evidence-based literature [36, 37], to reflect up-to-date, evidence-informed recommendations for enteral nutrition in critically ill patients. The policy is deterministic and constructed according to expert-derived nutritional targets and clinical rules. A full specification of this policy is provided in the Appendix (Section 3: Expert Guidelines Policy Definition).
@@ 3.2.4 Off-Policy Evaluation
    Average Returns: We used Consistent Weighted Per-Decision Importance Sampling (CWPDIS) [38] to estimate expected returns from each policy. As importance sampling based OPE methods require a behavior policy, we used our trained BC model baseline to estimate the behavior policy. A higher $V_{\text{CWPDIS}}$ corresponds to higher estimated returns and indicates a more effective policy.
@@ 4.2 Return-Mortality Correlation: Clinical Alignment
Figure 4: Dosage differences (x-axis) versus biomarker deviation (y-axis) for DeepEN. A) Glucose deviation vs calories dosage difference and B) Phosphate deviation vs protein dosage
@@
5 Discussion
DeepEN leverages deep reinforcement learning to dynamically personalize enteral nutrition (EN) strategies in critically ill patients, optimizing for hospital survival. The model is trained using an expert-informed, physiologically grounded reward function that integrates both long-term outcomes and short-term nutrition-specific signals. This design allows the model to assign lower value estimates to actions associated with higher mortality and higher value estimates to those linked with improved survival. In doing so, DeepEN learns a policy that aligns with high-value clinician actions while appropriately deviating from low-value ones. To ensure that recommendations remain clinically plausible, a conservative Q-learning penalty is incorporated to constrain the policy within the support of observed clinical practice.
@@ 5 Discussion
Although DeepEN’s observed mortality reduction of 3.7% (from 22.5% to 18.8%) is smaller than the improvements typically reported in reinforcement learning studies focused on fluid and vasopressor optimization[39, 40]—the most established and widely studied applications of healthcare RL—this effect remains clinically meaningful within the context of enteral nutrition therapy. The multifactorial nature of critical illness and the historically modest survival benefits demonstrated in large-scale nutrition trials[41, 3] underscore the significance of even incremental gains. Unlike hemodynamic interventions, which exert immediate effects on survival through cardiovascular stabilization, enteral nutrition confers its benefits more gradually through metabolic regulation, maintenance of gut integrity, and reduction of malnutrition-related complications. These mechanisms, though less acutely life-saving, are essential for supporting physiological recovery and long-term outcomes in critically ill patients.[1]
@@
While we train and evaluate DeepEN on a relatively large patient cohort, the data we use comes from a single hospital system. This may limit the generalizability of our model, as clinical practices and patient populations can vary substantially across institutions and regions. Validation on external cohorts will therefore be essential to confirm the robustness of our findings. However, the lack of detailed EN data in other large open-source critical care databases currently precludes this effort. We must also note that DeepEN’s reward function—while based heavily on well-established rewards functions in healthcare RL (cite)—is handcrafted. Consequently, it may not optimally capture all relevant aspects of nutritional adequacy or clinical recovery. The exploration and development of alternative rewards functions that better integrate and capture short- and long-term goals relevant to EN, represents an important future direction.
@@ 6 Ethical Statement
In line with prior work on reinforcement learning–driven clinical decision support systems (CDSSs) for conditions such as sepsis [11], our proposed framework adopts a human-in-the-loop approach. The system is designed to assist, rather than replace, the role of clinicians by offering data-driven recommendations derived from past clinical trajectories. Importantly, the ultimate responsibility for treatment decisions remains with the clinician, preserving professional autonomy and ensuring that medical expertise governs final care plans. This collaborative model enables the integration of machine intelligence with clinical judgment, leveraging the strengths of both to improve patient outcomes without undermining ethical standards.
@@ 3 Expert Guidelines Policy Definition
These nutritional targets are based on the 2021 ASPEN requirements for enteral nutrition [3]. Original protein targets have been altered slightly with more recent evidence-based literature recommending a low-dosing period during the early-acute phase of critical illness [36, 37].
@@ 3.3 Water
  • 1.5 ml per 1 kcal of calories administered.

### v2

@@ 3.2.2 Baselines
    Expert Guidelines (EG) Policy: This rule-based policy encodes clinical recommendations drawn from the American Society for Parenteral and Enteral Nutrition (ASPEN) guidelines [compher2022guidelines], in combination with more recent evidence-based literature [wischmeyer2023personalized, stoian2025personalized], to reflect up-to-date, evidence-informed recommendations for enteral nutrition in critically ill patients. The policy is deterministic and constructed according to expert-derived nutritional targets and clinical rules. A full specification of this policy is provided in the Appendix (Section 3: Expert Guidelines Policy Definition).
@@ 3.2.4 Off-Policy Evaluation
    Average Returns: We used Consistent Weighted Per-Decision Importance Sampling (CWPDIS) [thomas2016data] to estimate expected returns from each policy. As importance sampling based OPE methods require a behavior policy, we used our trained BC model baseline to estimate the behavior policy. A higher $V_{\text{CWPDIS}}$ corresponds to higher estimated returns and indicates a more effective policy.
@@ 4.2 Return-Mortality Correlation: Clinical Alignment
Figure 4: Dosage differences (x-axis) versus biomarker deviation (y-axis) for DeepEN. (A) Glucose deviation vs. calories dosage difference and (B) Phosphate deviation vs. protein dosage.
@@
5 Discussion and Conclusions
DeepEN applies deep RL to dynamically personalize enteral nutrition for critically ill patients, with the overarching goal of improving hospital survival. The framework is trained using an expert-informed state space and a physiologically grounded reward function that integrates long-term outcomes with short-term nutrition-relevant signals. Together, these design choices enable DeepEN to effectively distinguish actions associated with higher mortality from those linked to better survival, producing a policy that aligns with high-value clinician decisions while appropriately deviating from suboptimal ones. To ensure clinical plausibility, the framework incorporates a conservative Q-learning penalty that constrains the learned policy to remain within the support of observed clinical practice.
@@ 5 Discussion and Conclusions
Although DeepEN’s observed mortality reduction of 3.7% (from 22.5% to 18.8%) is smaller than the improvements typically reported in healthcare RL studies focused on fluid and vasopressor dosage optimization [wu2023value, komorowski2018artificial], this effect remains clinically meaningful within the context of enteral nutrition therapy. The multifactorial nature of critical illness and the historically modest survival benefits demonstrated in large-scale nutrition trials [bear2017role, compher2022guidelines] underscore the significance of even incremental gains. Unlike hemodynamic interventions, which exert immediate effects on survival through cardiovascular stabilization, enteral nutrition confers its benefits more gradually through metabolic regulation, maintenance of gut integrity, and reduction of malnutrition-related complications. These mechanisms, though less acutely life-saving, are essential for supporting physiological recovery and long-term outcomes in critically ill patients. [adeyinka2018enteric]
@@
While we train and evaluate DeepEN on a relatively large patient cohort, the data we use comes from a single hospital system. This may limit the generalizability of our model, as clinical practices and patient populations can vary substantially across institutions and regions. Validation on external cohorts will therefore be essential to confirm the robustness of our findings. However, the lack of detailed EN data in other large open-source critical care databases currently precludes this effort. We must also note that DeepEN’s reward function—while based heavily on well-established rewards functions in healthcare RL [raghu2017deep, tan2024advancing, wu2023value]—is handcrafted. Consequently, it may not optimally capture all relevant aspects of nutritional adequacy or clinical recovery. The exploration and development of alternative rewards functions that better integrate and capture short- and long-term goals relevant to EN, represents an important future direction.
@@ 6 Ethical Statement
In line with prior work on reinforcement learning–driven clinical decision support systems (CDSSs) for conditions such as sepsis [raghu2017deep], our proposed framework adopts a human-in-the-loop approach. The system is designed to assist, rather than replace, the role of clinicians by offering data-driven recommendations derived from past clinical trajectories. Importantly, the ultimate responsibility for treatment decisions remains with the clinician, preserving professional autonomy and ensuring that medical expertise governs final care plans. Our framework enables the integration of machine intelligence with clinical judgment, leveraging the strengths of both to improve patient outcomes without undermining ethical standards.
@@ 8 Code Availability
Acknowledgment
The authors would like to acknowledge the contributions of their team members and mentors from the Singapore IMAGINE AI 2024 Datathon.
@@ 3 Expert Guidelines Policy Definition
These nutritional targets are based on the 2021 ASPEN requirements for enteral nutrition [compher2022guidelines]. Original protein targets have been altered slightly with more recent evidence-based literature recommending a low-dosing period during the early-acute phase of critical illness [wischmeyer2023personalized, stoian2025personalized].
@@ 3.3 Water
  • 1.
    1.5 ml per 1 kcal of calories administered.